In H.C Vermas concepts of physics vol-2 in the page.no.106 in Example 29.1, what is the rule behind taking E=Eacos@ + Ebcos@ ????


Why cant sin@ be taken instead of cos@ ???


 


Pls...help Pls i will approve the ans anyway help me aieee approaching n givin me the spine chills !!!


 


Vikas waiting 4 ur reply !!! i know u knw this come on help me pls.....

Grade:12

2 Answers

vikas askiitian expert
13 years ago

electric field is a vector quantity so it can be broken into components ,

 

Ea is electric field by charge which is placed at vertex A

Eb is electric field by charge which is placed at vertex B

 

now , Ea & Eb both can be divided into two components ,

for Ea two components are Eacos@ & Easin@

for Eb two components are Ebcos@ & Ebsin@

 

now , it is clear that Eb sin@ is in opposite direction to Easin@ so they will cancel the effect of each other but

on other side Eacos@ & Ebcos@ are acting olong same direction so they will be add up to give total

electric field E = Eacos@ + Ebcos@

 

 

instead of all these one other method can be used for this particular problem , see both charges are of same magnitude

but opposite sign so this system will behave as a dipole & for this we can use direct formula

E(equitorial) = kp/(a2+x2)3/2

here p(dipole moment) = q(2a)

     x = perpendicular distance

     2a is distance bw charges

XYZ Person
13 years ago

drop a pependicular from PEa to PE now in triangle PEEa

cos@=E/Ea

=> Ea=E cos@

 similarly Eb=E cos@

there is no rule behind its just d simple trigonometry

 

tejaswi